# Multi-color XFEL pulses with variable color separation and time delay for multi-frame diffraction imaging

### v1

@@ Abstract
X-ray free-electron lasers (XFELs) of high brightness have opened new opportunities for exploring ultrafast dynamical processes in matter, enabling imaging and movies of single molecules and particles at atomic resolution. In this paper, we present a straightforward method for multi-frame diffraction imaging, using the same electron beam to generate four-color XFEL pulses with adjustable wavelength separation and time delay. The optical klystron scheme is introduced to enhance FEL intensity and reduce the total length of undulators. The time delay is tuned via a magnetic chicane between the undulators with various colors. Using parameters of SHINE, start-to-end simulations demonstrate the effectiveness and tunability of our method, achieving representative results such as time delays of hundreds of femtoseconds and four-color XFEL pulses spanning 1.8 to 2.7 nm with 0.3 nm intervals. The proposed scheme enables the recording of multi-frame diffraction images in a single exposure, providing a new perspective for ultrafast molecular and atomic dynamics studies.
keywords:
multi-color; free-electron lasers; optical klystron; multi-frame diffraction imaging
1 Introduction
@@
In recent years, X-ray free-electron lasers (XFELs)1, 2, 3 have provided revolutionary tools for exploring ultrafast dynamics in matter due to their ultrashort pulses (femtosecond to attosecond durations), high brightness, and coherence. In time-resolved studies, pump-probe configurations that employ either a conventional optical laser or an FEL-generated X-ray pump combined with an X-ray FEL probe are indispensable for capturing ultrafast dynamics across physical phenomena 4, 5, 6, 7, 8, 9, 10, chemical reactions 11, 12, 13, and biological structures 14, 15 on the picosecond to femtosecond time scale. However, these methods rely on repetitive measurements for image construction, which precludes their application to non-repetitive or irreproducible events 16. To fully track ultrafast dynamical processes, it is necessary to continuously record multiple transient moments in a single experiment, akin to high-speed photography or flash imaging 17. Thus, multi-color XFEL pulses with controllable time delays are urgently required.
@@
Various multi-color XFEL schemes have been proposed by manipulating electron beam properties, including twin bunches 18, two-bucket bunches 19, and the use of different energy parts within one bunch 20, double-slot foil 21, sextupole 22, nonlinear electron compression 23 and laser emittance spoiler 24. Meanwhile, artificially changing the undulator parameter can straightforwardly produce tunable multi-color XFEL pulses 25, 26, 27. However, when the whole electron beam is utilized to generate multi-color XFEL pulses, the downstream color intensity is constrained by the upstream color induced energy spread. The fresh-slice technique enables precise slicing control of the electron beam, allowing the fresh portion to undergo independent lasing in different undulators, thereby achieving high power and multi-color XFEL pulses 28, 29, 30, 31, 32.
@@
Unlike the fresh-slice technique, which requires additional electron beam matching and orbit optimization, in this paper, we propose a multi-frame X-ray diffraction imaging scheme in which four-color XFEL pulses are generated by the whole electron bunch. Considering the limitation of XFEL-induced energy spread and the undulator length, the optical klystron (OK) scheme is introduced to speed up XFEL growth while shortening undulator length requirements 33, 34, 35, 36, 37, 38, which enables the generation of high-power multi-color XFEL pulses and offers straightforward implementation in existing XFEL facilities. The proposed scheme can generate four discrete-wavelength XFEL pulses with tunable wavelength separation and time delay. The pulse energy of each color can be adjusted using the $R_{56}$ of the dispersive chicane. Moreover, the feasibility of the proposed scheme is confirmed by the reflection grating diffraction results. The tailored XFEL pulses sequentially expose the sample to capture temporal dynamics, with the diffracted signals spatially dispersed by a grating and simultaneously recorded as four distinct diffraction patterns using an X-ray CCD detector. This scheme can leverage the unique properties of multi-color XFEL pulses, enabling time-resolved flash imaging and facilitating ultrafast dynamical studies of matter with high temporal resolution.
@@
This paper is organized as follows. Section 2 elaborates on the principles of the proposed scheme. Numerical simulations of multi-color XFEL generation and transport based on the main parameters of Shanghai High Repetition Rate XFEL and Extreme Light Facility (SHINE) are presented in Section 3 and  4. Conclusion and discussion are described in Section 5.
2 Principles
@@
The proposed scheme can be divided into two phases, as shown in Fig. 1. First, the generation of multi-color XFEL pulses employs the electron beam passing through these undulator modules ($A_{N}$, $B_{N}$, $C_{N}$, and $D_{N}$, each with distinct undulator parameters $K$), and three time-delay chicanes. As illustrated in Fig. 1(a), each undulator module consists of a sub-undulator, a small dispersive chicane, and a quadrupole. Second, the high-time-resolution dynamic diffraction imaging system features the four-color X-ray pulses interacting with a sample, followed by a high line-density reflective grating, and an X-ray CCD detector, as depicted in Fig. 1(b).
@@
Figure 1: Proposed multi-frame X-ray diffraction imaging scheme. (a) Schematic of the multi-color XFEL pulses generation. Components include an electron beam (bright yellow), sub-undulators (blue-and-white striped), dispersive chicanes (purple), quadrupole magnets (green-and-red striped), time-delayed chicanes (gray), and an electron beam collector (black). Four-color XFEL pulses, each with distinct wavelengths, are represented by unique colors (color-coded). (b) Schematic of the high-time-resolution dynamic diffraction imaging system. This includes a sample (gray sphere), a high line-density reflective grating (bright yellow), and an X-ray CCD detector (black). The four X-ray pulses sequentially illuminate the sample to capture temporal information of ultrafast events, which are dispersed spatially via the grating and recorded as four diffraction images by the CCD detector.
@@
As the electron beam propagates through sub-undulator $A_{1}$, FEL radiation is generated at the resonance wavelength, given by 1
@@
where $\lambda_{u}$ denotes the undulator period, $K$ represents the undulator parameter, and $\gamma$ is the Lorentz factor. In the amplification process of the OK-enhanced self-amplified spontaneous emission (SASE) scheme 37, the energy modulation is converted to density modulation via a small dispersive chicane. Moreover, this configuration speeds the electron beam’s microbunching, thereby reducing the FEL gain length. Notably, the OK efficiency is substantially enhanced when the electron beam’s rms relative energy spread $\sigma_{\delta}$ is much smaller than the FEL Pierce parameter $\rho$ 1. The one-dimensional theory describing the maximum theoretical power gain factor in the OK-SASE occurs when 39:
@@
where $R_{56}$ denotes the momentum compaction of the small dispersive chicane. However, if the first sub-undulator $A_{1}$ is sufficiently long, it can generate an energy modulation with an amplitude that is comparable to or exceeds the intrinsic energy spread. In this scenario, Equation (2) no longer corresponds to the maximum gain, as the induced energy modulation becomes dominant over the intrinsic energy spread. By considering the high-gain harmonic generation 40, we can analyze the scenario that may arise in the aforementioned situation 37. The nth harmonic bunching factor can be derived as:
@@
where $J_{n}$ is the $n$th order Bessel function, $\Delta E$ denotes the energy modulation amplitude induced by the first sub-undulator $A_{1}$ and $E$ represents the mean energy of the electron bunch. From Equation (3), we can analyze the OK behavior of $R_{56}$ to the effects of different energy modulation amplitudes, specifically $\Delta E$ = 250 keV and $\Delta E$ = 2000 keV, for an electron bunch with a mean energy of $E$ = 4.48 GeV and an intrinsic energy spread of 650 keV ($\sigma_{\delta}=1.45\times 10^{-4}$). Furthermore, undulator modules ($A_{1}-A_{N}$) are all tuned to resonate at a wavelength of 1.8 nm, and only the fundamental harmonic ($n=1$) is considered. Figure 2(a) illustrates the scenario where the relative energy modulation amplitude is smaller than the rms relative intrinsic energy spread, i.e., $\frac{\Delta E}{E}<\sigma_{\delta}$, the contribution of the exponential decay to the bunching factor becomes more significant than that of the Bessel function. In this case, the value of $R_{56}^{1}$ corresponding to the maximum bunching is found to be $1.94\ \mu m$. Notably, this value matches optimal $R_{56}^{1}$ calculated from Eq. (2), which also yields $1.97\ \mu m$. This agreement confirms that the bunching factor reaches its maximum when the condition described by Eq. (2) is satisfied. In contrast, as shown in Fig. 2(b), when the Bessel function dominates, the optimal value of $R_{56}^{1}$ shifts to $1\ \mu m$, where the maximum bunching factor is achieved under the condition $\frac{\Delta E}{E}k_{r}R_{56}\approx 1.8412$. This highlights the transition between the dominance of the exponential decay and the Bessel function in determining the bunching behavior, with each regime yielding distinct optimal values of $R_{56}^{1}$ corresponding to their respective conditions for maximizing the bunching factor.
@@
When the bunched electron beam passes through the sub-undulator $A_{N-1}$, it undergoes N2N-2 stages of OK-enhanced SASE FEL amplification, specifically referred to as the seeded FEL amplification. During this process, the energy modulation amplitude induced by the sub-undulator $A_{N-1}$ before the last dispersive chicane $R_{56}^{N-1}$ becomes larger than the intrinsic energy spread. As the energy modulation amplitude increases, the optimal value of $R_{56}^{N-1}$ decreases due to the dominance of the Bessel function governing the bunching process. Specifically, a larger energy modulation amplitude requires a smaller value $R_{56}^{N-1}$ to maximize the bunching factor, as shown in Fig. 3. This optimization ultimately enhances FEL radiation in the subsequent undulator module $A_{N}$. The time delay of delay chicane 1 is twice its $R_{56}$, which is sufficient to smear out the microbunching generated by the previous undulator module $A_{N}$. For instance, when $R_{56}=600\ \mu m$, the bunching factor is nearly zero, as given by Eq. (3). Additionally, the delay ensures that the electron beam is properly delayed, allowing downstream modules to re-establish the FEL process. Similarly, the multi-stage OK configuration involves (N1N-1) small dispersive chicanes for each wavelength, including $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$.
@@
Therefore, four-color XFEL pulses with distinct wavelengths ($\lambda_{1}<\lambda_{2}<\lambda_{3}<\lambda_{4}$) and temporal intervals between adjacent pulses ranging from picoseconds to femtoseconds provide a novel experimental methodology for ultrafast diffraction experiments utilizing multiple monochromatic short pulses in Fig. 1(b). When a set of four-color pulses train within a single exposure is generated and sequentially interacts with an object, and the pulses are diffracted by a grating with high line-density. At the grating’s image plane, diffracted light from pulses of different colors ultimately falls at distinct positions. Consequently, information carried by different pulses within the set is recorded at separate image plane locations. Furthermore, the temporal delay between pulses enables a four-color pulse train to comprehensively capture material information over a specific time interval. Upon passing through a high line-density reflective grating 41, the X-rays can be spatially separated according to the grating equation:
@@
the resolution power of the grating is determined by the relationship $R=\frac{\lambda}{\Delta\lambda_{min}}=mN$, where $\Delta\lambda_{min}$ denotes the minimum resolvable wavelength difference and $N$ is the total number of illuminated grooves on the grating, given by $N=\frac{L}{d}$, with $L$ being the projected length of the beam on the grating surface. For the four-color XFEL pulses, the FWHM bandwidth of each wavelength was greater than twice its specific minimum resolvable wavelength difference, and the wavelength difference between adjacent pulses is sufficiently large such that it exceeds the minimum resolvable wavelength differences. As a result, the X-rays are clearly separated by the grating, spatially dispersed in ascending order of wavelength ($\lambda_{1}<\lambda_{2}<\lambda_{3}<\lambda_{4}$), and symmetrically arranged around the zeroth-order diffraction peak. Each X-ray pulse generates a unique diffraction angle via the grating, resulting in angles $\theta_{1}^{\prime}$, $\theta_{2}^{\prime}$, $\theta_{3}^{\prime}$, and $\theta_{4}^{\prime}$, allowing for clear and distinct separation of the four X-ray pulses.
@@
In the proposed scheme, adjusting the time delay parameter of the delay chicane controls the inter-pulse time separation between adjacent X-ray pulses to set the frame interval. Tuning the $K$ parameter of sub-undulators regulates both the wavelength separation and specific X-ray wavelengths. The energy of each X-ray pulse can be effectively controlled by jointly tuning the number of undulator modules and the parameters $R_{56}$ of dispersive chicanes, where $R_{56}$ is optimized by the differential evolution algorithm – a stochastic population-based algorithm employing vector differential mutation, binomial crossover, and competitive selection mechanisms 42. Theoretically, this scheme enables the generation of multi-color X-ray pulses for multi-frame diffraction imaging. However, limitations in adjustable frame numbers exist due to electron beam energy spread growth and the dynamic range constraints of the X-ray CCD, which are further influenced by its physical size.
3 Start-to-end XFEL simulation
In this section, we explore the generation of multi-color XFEL pulses at the soft X-ray beamline of SHINE 43. The SHINE employs an 8 GeV CW superconducting RF linac 44, 45 to deliver X-ray photons from 0.4-25 keV at a 1 MHz repetition rate, with two FEL beamlines: hard X-ray (FEL-I) and soft X-ray (FEL-II). Besides 8 GeV full beam energy operation, FEL-II beamline 43 can operate with a 3.6-4.5 GeV electron beam and support various modes, including SASE, self-seeding, and external seeded FEL schemes at MHz repetition rate46, 47. Its undulator system, featuring a 55 mm period length, comprises 32 undulator modules with each 4 m long. Potentially, small dispersive chicanes can be integrated within the 1 m interspaces between these modules, thereby forming reconfigured functional modules.
The start-to-end (S2E) simulation was performed using ASTRA48 for the electron beam dynamics in the photoinjector, ELEGANT49 for the linac transport including coherent synchrotron radiation effects, and Genesis 1.3 50 for the three-dimensional FEL simulations. The electron beam, whose current profile and longitudinal phase space properties are shown in Fig. 4, is characterized by the parameters listed in Table 1.

### v2

# Multi-color x-ray free-electron laser generation using optical klystron for multi-frame diffraction imaging

@@ Abstract
X-ray free-electron lasers (XFELs) of high brightness have opened new opportunities for exploring ultrafast dynamical processes in matter, enabling imaging and movies of single molecules and particles at atomic resolution. In this paper, we present a straightforward method for multi-frame diffraction imaging, using the whole electron beam to generate four-color XFEL pulses with adjustable wavelength separation and time delay. The optical klystron scheme is introduced to enhance FEL intensity and reduce the total length of undulators. The time delay is tuned via a magnetic chicane between the undulators with various colors. Using parameters of SHINE, start-to-end simulations demonstrate the effectiveness and tunability of our method, achieving representative results such as time delays of hundreds of femtoseconds and four-color XFEL pulses spanning 1.8 to 2.7 nm with 0.3 nm intervals. The proposed scheme enables the recording of multi-frame diffraction images in a single exposure, providing a new perspective for ultrafast molecular and atomic dynamics studies.
preprint: APS/123-QED
I Introduction
@@
In recent years, X-ray free-electron lasers (XFELs) [23, 33, 22] have emerged as transformative tools for probing ultrafast dynamics in matter, owing to their femtosecond-to-attosecond pulse durations, high brightness, and full transverse coherence. In time-resolved studies, pump-probe configurations that employ either a conventional optical laser or an FEL-generated X-ray pump combined with an X-ray FEL probe are indispensable for capturing ultrafast dynamics across physical phenomena [36, 1, 24, 18, 15, 37, 9, 21], chemical reactions [2, 30, 3], and biological structures [6, 49] on the picosecond to femtosecond time scale. However, these methods rely on shot-to-shot repetitive measurements with scanned pump-probe delays for image construction, which limits their application to non-repetitive, irreversible, or poorly reproducible events [31]. In such cases, it is highly desirable to record multiple transient states originating from the same excitation event in a single shot, analogous to high-speed photography or flash imaging [46]. Therefore, single-shot multi-color XFEL pulses with well-defined and controllable relative time delays provide an effective approach for enabling multi-frame diffraction imaging of ultrafast dynamics.
@@
Various multi-color XFEL schemes have been proposed by manipulating electron beam properties, including twin bunches [29], two-bucket bunches [11], and the use of different energy parts within one bunch [44], double-slot foil [45], sextupole [13], nonlinear electron compression [28] and laser emittance spoiler [48]. Meanwhile, artificially changing the undulator parameter can straightforwardly produce tunable multi-color XFEL pulses [26, 20, 10]. However, when the whole electron bunch is used to generate multiple colors sequentially, the intensity of downstream pulses is limited by the energy spread induced by upstream lasing. The fresh-slice technique circumvents this limitation by selectively utilizing unmodulated portions of the beam for independent amplification in separate undulators, thereby enabling high-power multi-color operation [27, 43, 19, 7, 39].
@@
Inspired by a single-shot multi-frame imaging concept previously demonstrated at optical wavelengths [46], we propose an extension to XFELs. In contrast to fresh-slice schemes, the proposed scheme generates four-color XFEL pulses from the whole electron bunch, enabling single-shot multi-frame X-ray diffraction imaging with adjustable wavelength separation and time delay. To mitigate the constraints imposed by XFEL-induced energy spread and undulator length, we employ an optical klystron (OK) scheme [4, 32, 47, 34, 35, 38], which enhances gain and reduces saturation length, facilitating high-power multi-color pulse generation with straightforward implementation in existing facilities. The proposed scheme can generate four discrete-wavelength XFEL pulses with tunable wavelength separation and time delay, and the pulse energy of each color can be adjusted using the $R_{56}$ of the dispersive chicane. Moreover, the feasibility of the proposed scheme is confirmed by the reflection grating diffraction results. In operation, a single pump pulse initiates a reaction, followed by three or four XFEL probe pulses with femtosecond- to picosecond-scale delays. These tailored XFEL pulses sequentially probe the same reaction. The resulting diffraction signals are spatially separated by a grating and simultaneously recorded on an X-ray CCD detector. Since all pulses originate from the same electron bunch, they are intrinsically synchronized, which reduces relative timing uncertainties and thereby enables single-shot, multi-frame time-resolved flash imaging of ultrafast dynamics.
@@
This paper is organized as follows. Section II elaborates on the principles of the proposed scheme. Numerical simulations of multi-color XFEL generation and transport based on the main parameters of Shanghai High Repetition Rate XFEL and Extreme Light Facility (SHINE) are presented in Section III and  IV. Conclusion and discussion are described in Section V.
II Principles
@@
The proposed scheme can be divided into two phases, as shown in Fig. 1. First, the generation of multi-color XFEL pulses employs the electron beam passing through these undulator modules ($A_{N}-D_{N}$, each with distinct undulator parameters $K$), and three time-delay chicanes as shown in Fig. 1(a) and Fig. 1(b). Second, the high-time-resolution dynamic diffraction imaging system features the four-color X-ray pulses interacting with a sample, followed by a high line-density reflective grating, and an X-ray CCD detector, as depicted in Fig. 1(c).
@@
Figure 1: Proposed multi-frame X-ray diffraction imaging scheme. (a) Schematic layout of the multi-color XFEL pulses generation. The setup consists of an electron beam, four undulator modules denoted as $A_{N}$, $B_{N}$, $C_{N}$ and $D_{N}$, three time-delayed chicanes, and an electron beam collector, where N denotes the number of sub-undulators in each module. The modules generate XFEL pulses at wavelengths $\lambda_{1}-\lambda_{4}$, color-coded by wavelength. The dashed box indicates an undulator module detailed in panel (b), which comprises N sub-undulators $U_{X,i}\ (X=A,B,C,D;\ i=1,\dots,N)$, interleaved with N-1 small dispersive chicanes. (c) Schematic layout of the high-time-resolution dynamic diffraction imaging system. This includes a sample, a high line-density reflective grating, and an X-ray CCD detector. The four X-ray pulses sequentially illuminate the sample to capture temporal information of ultrafast events, which are dispersed spatially via the grating and recorded as four diffraction images by the CCD detector.
@@
As the electron beam propagates through sub-undulator $U_{A,1}$, FEL radiation is generated at the resonance wavelength, given by [23]
@@
where $\lambda_{u}$ denotes the undulator period, $K$ represents the undulator parameter, and $\gamma$ is the Lorentz factor. In the amplification process of the OK-enhanced self-amplified spontaneous emission (SASE) scheme [35], the energy modulation is converted to density modulation via a small dispersive chicane. Moreover, this configuration speeds the electron beam’s microbunching, thereby reducing the FEL gain length. Notably, the OK efficiency is substantially enhanced when the electron beam’s rms relative energy spread $\sigma_{\delta}$ is much smaller than the FEL Pierce parameter $\rho$ [23]. The one-dimensional theory describing the maximum theoretical power gain factor in the OK-SASE occurs when [14]:
@@
where $R_{56}$ denotes the momentum compaction of the small dispersive chicane. However, if the first sub-undulator $U_{A,1}$ is sufficiently long, it can generate an energy modulation with an amplitude that is comparable to or exceeds the intrinsic energy spread. In this scenario, Equation (2) no longer corresponds to the maximum gain, as the induced energy modulation becomes dominant over the intrinsic energy spread. By considering the high-gain harmonic generation [52], we can analyze the scenario that may arise in the aforementioned situation [35]. The nth harmonic bunching factor can be derived as:
@@
where $J_{n}$ is the $n$th order Bessel function, $\Delta E$ denotes the energy modulation amplitude induced by the first sub-undulator $U_{A,1}$ and $E$ represents the mean energy of the electron bunch. From Equation (3), we can analyze the OK behavior of $R_{56}$ to the effects of different energy modulation amplitudes, specifically $\Delta E$ = 250 keV and $\Delta E$ = 2000 keV, for an electron bunch with a mean energy of $E$ = 4.5 GeV and an intrinsic energy spread of 650 keV ($\sigma_{\delta}=1.44\times 10^{-4}$). Furthermore, the undulator module $A_{N}$ is tuned to resonate at a wavelength of 1.8 nm, and only the fundamental harmonic ($n=1$) is considered. Figure 2(a) illustrates the scenario where the relative energy modulation amplitude is smaller than the rms relative intrinsic energy spread, i.e., $\frac{\Delta E}{E}<\sigma_{\delta}$, the contribution of the exponential decay to the bunching factor becomes more significant than that of the Bessel function. In this case, the value of $R_{56}^{1}$ corresponding to the maximum bunching is found to be $1.94\ \mu m$. Notably, this value matches optimal $R_{56}^{1}$ calculated from Eq. (2), which also yields $1.98\ \mu m$. This agreement confirms that the bunching factor reaches its maximum when the condition described by Eq. (2) is satisfied. In contrast, as shown in Fig. 2(b), when the Bessel function dominates, the optimal value of $R_{56}^{1}$ shifts to $1\ \mu m$, where the maximum bunching factor is achieved under the condition $\frac{\Delta E}{E}k_{r}R_{56}\approx 1.8412$. This highlights the transition between the dominance of the exponential decay and the Bessel function in determining the bunching behavior, with each regime yielding distinct optimal values of $R_{56}^{1}$ corresponding to their respective conditions for maximizing the bunching factor.
@@
When the bunched electron beam passes through the sub-undulator $U_{A,N-1}$, it undergoes N2N-2 stages of OK-enhanced SASE FEL amplification, specifically referred to as the seeded FEL amplification. During this process, the energy modulation amplitude induced by the sub-undulator $U_{A,N-1}$ before the last dispersive chicane $R_{56}^{N-1}$ becomes larger than the intrinsic energy spread. As the energy modulation amplitude increases, the optimal value of $R_{56}^{N-1}$ decreases due to the dominance of the Bessel function governing the bunching process. Specifically, a larger energy modulation amplitude requires a smaller value $R_{56}^{N-1}$ to maximize the bunching factor, as shown in Fig. 3. This optimization ultimately enhances FEL radiation in the subsequent sub-undulator $U_{A,N}$. The time delay of delay chicane 1 is twice its $R_{56}$, which is sufficient to smear out the microbunching generated by the previous undulator module $A_{N}$. For instance, when $R_{56}=600\ \mu m$, the bunching factor is nearly zero, as given by Eq. (3). Similarly, the multi-stage OK configuration involves N1N-1 small dispersive chicanes for each wavelength, including $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$.
@@
Therefore, four-color XFEL pulses with distinct wavelengths ($\lambda_{1}<\lambda_{2}<\lambda_{3}<\lambda_{4}$) and temporal intervals between adjacent pulses ranging from picoseconds to femtoseconds provide a novel experimental methodology for ultrafast diffraction experiments utilizing multiple monochromatic short pulses in Fig. 1(c). When a set of four-color pulses is generated within a single exposure and sequentially interacts with an object, and the pulses are diffracted by a grating with high line-density. At the grating’s image plane, diffracted light from pulses of different colors ultimately falls at distinct positions. Consequently, information carried by different pulses within the set is recorded at separate image plane locations. Furthermore, the temporal delay between pulses enables a four-color pulse train to comprehensively capture material information over a specific time interval. Upon passing through a high line-density reflective grating [41], the X-rays can be spatially separated according to the grating equation:
@@
the resolving power of the grating is determined by the relationship $R=\frac{\lambda}{\Delta\lambda_{min}}=mN$, where $\Delta\lambda_{min}$ denotes the minimum resolvable wavelength difference and $N$ is the total number of illuminated grooves on the grating, given by $N=\frac{L}{d}$, with $L$ being the projected length of the beam on the grating surface. For the four-color XFEL pulses, the full width at half maximum (FWHM) bandwidth of each wavelength was greater than twice its specific minimum resolvable wavelength difference, and the wavelength difference between adjacent pulses is sufficiently large such that it exceeds the minimum resolvable wavelength differences. As a result, the X-rays are clearly separated by the grating, spatially dispersed in ascending order of wavelength ($\lambda_{1}<\lambda_{2}<\lambda_{3}<\lambda_{4}$), and symmetrically arranged around the zeroth-order diffraction peak. Each X-ray pulse generates a unique diffraction angle via the grating, resulting in angles $\theta_{1}^{\prime}$, $\theta_{2}^{\prime}$, $\theta_{3}^{\prime}$, and $\theta_{4}^{\prime}$, allowing for clear and distinct separation of the four X-ray pulses.
@@
In the proposed scheme, adjusting the time delay of the large chicane controls the inter-pulse time separation between adjacent X-ray pulses to set the frame interval. Tuning the $K$ parameter of sub-undulators regulates both the wavelength separation and specific X-ray wavelengths. The energy of each X-ray pulse can be effectively controlled by jointly tuning the number of sub-undulators and the parameters $R_{56}$ of dispersive chicanes, where $R_{56}$ is optimized by the differential evolution algorithm – a stochastic population-based algorithm employing vector differential mutation, binomial crossover, and competitive selection mechanisms [53]. Theoretically, this scheme enables the generation of multi-color X-ray pulses for multi-frame diffraction imaging. However, limitations in adjustable frame numbers exist due to electron beam energy spread growth and the dynamic range constraints of the X-ray CCD, which are further influenced by its physical size.
III Start-to-end XFEL simulation
In this section, we explore the generation of multi-color XFEL pulses at the soft X-ray beamline of SHINE [25]. The SHINE employs an 8 GeV CW superconducting RF linac [50, 8] to deliver X-ray photons from 0.4-25 keV at a 1 MHz repetition rate, with two FEL beamlines: hard X-ray (FEL-I) and soft X-ray (FEL-II). Besides 8 GeV full beam energy operation, FEL-II beamline [25] can operate with a 3-4.5 GeV electron beam and support various modes, including SASE, self-seeding, and external seeded FEL schemes at MHz repetition rate [51, 40]. Its undulator system, featuring a 55 mm period length, comprises 32 undulator modules, each 4 m long. Potentially, small dispersive chicanes can be integrated within the 1 m interspaces between these modules, thereby forming reconfigured functional modules.
The start-to-end (S2E) simulation was performed using ASTRA [16] for the electron beam dynamics in the photoinjector, ELEGANT [5] for the linac transport including coherent synchrotron radiation effects, and Genesis 1.3 [42] for the three-dimensional FEL simulations. The electron beam, whose current profile and longitudinal phase space properties are shown in Fig. 4, is characterized by the parameters listed in Table 1. In the multi-color FEL simulations, the particle distribution is transferred between modules, and wavelength switching is handled by reslicing the beam according to the new resonant wavelength.